Solve the system, if possible. (If the system is inconsistent, enter INCONSISTENT. If the system is dependent, enter DEPENDENT.)
2x + 2y + 3z = 20
3x + y - z = 0
x + y + 2z = 12

Answer by TimothyLamb(4379)

---
2x + 2y + 3z = 20
3x + y - z = 0
x + y + 2z = 12
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 0
y = 4
z = 4
---
the system is INDEPENDENT.
---
system matrix
2 2 3
3 1 -1
1 1 2
---
inverse of system matrix
-1.5 0.5 2.5
3.5 -0.5 -5.5
-1 0 2
---
determinant of system matrix = -2
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php